OpenStudy (anonymous):

Let f(x) = (-3x^2 +16x +c) / (x^3 - 5x^2) Determine teh constant c such that lim x->5 f(x) exists.

OpenStudy (amistre64):

you want to make a hole here so that the graph jumps over the spot; as opposed to creating a vertical asymptote i believe

OpenStudy (amistre64):

T = -3x^2 +16x +c; (-3x-B)(x-5) D = x^3 - 5x^2; x^2 (x-5)

OpenStudy (amistre64):

does that make sense? we are trying to control the offending denominator with its likeness up top

OpenStudy (anonymous):

um what does t and d stand for?

OpenStudy (amistre64):

"t"op and "d"enominator .... just names

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

what we need to do is get it so (x-5) is a factor of the top part ...

OpenStudy (amistre64):

-3x^2 +16x +c = (-3x-B) (x-5) -3x-B x-5 ------ -3x^2 -Bx +15x +15 --------------- -3x^2+16x +15 B=-1

OpenStudy (amistre64):

..... let me try that again

OpenStudy (amistre64):

-3x-B x-5 ------ -3x^2 -Bx +15x +5B --------------- -3x^2+16x +5B B=-1; c=-5 is what i think i meant to say .. lets try it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im not sure i understnad what this means

OpenStudy (amistre64):

it means that the limit from the left and the right are the same when x approaches 5.

OpenStudy (amistre64):

if we were not able to control that offending part of the denominator; there would be no limit at x=5

OpenStudy (anonymous):

but how do we show work for this??

OpenStudy (amistre64):

factor the bottom to see the bad part; create it as a factor for the top; and solve

OpenStudy (anonymous):

wait is teh wrk what u showed b4

OpenStudy (amistre64):

yes, i try to work it the stated problem as a means to solve it ... lol

OpenStudy (amistre64):

\[\lim_{x->5}\frac{-3x+16x+c}{x^3-5x^2}\] \[\lim_{x->5}\frac{-3x+16x+c}{x^2(x-5)}\] \[\lim_{x->5}\frac{(Ax+B)(x-5)}{x^2(x-5)}\] where (Ax+B)(x-5) = -3x+16x+c

OpenStudy (amistre64):

\[(Ax+B)(x-5) = -3x+16x+c\] \[Ax^2+(-5A+B)x-5B= -3x+16x+c\] and equate coefficients \[Ax^2=-3x^2\]\[(-5A+B)x=16x\]\[-5B=c\] \[A=-3\]\[-5(-3)+B=16;\ B=1\]\[-5(1)=c\]

OpenStudy (anonymous):

wow i wud have never thought of this solution by myself. This makes so much more sense. If i have to find a constat but tehre is no denominaotr what shud i do?

OpenStudy (amistre64):

if there is no denominator ..... well, other than it just being a polynomial, id have to have an example of what you mean.

OpenStudy (amistre64):

it is because of the denominator here that we get a bad value at x=5

OpenStudy (amistre64):

polynomials as a rule are continuous and need no control

OpenStudy (anonymous):

here is an example: If g(x) = x^3 - x^2 + x , show taht a there exist a c at g(c) =10

OpenStudy (amistre64):

that is given a polynomial and has no bad parts; you just have to solve it for f(x) = 10.

OpenStudy (amistre64):

err... g(x) in this case

OpenStudy (amistre64):

it aint pretty in the end, but it is real

OpenStudy (anonymous):

is that supposed to b teh answer??

OpenStudy (anonymous):

it looks waaay 2 confusing XD

OpenStudy (amistre64):

i dont think that is the answer. but it helps to see that an answer exists. now as i see it, they are not asking you to locate c, but to show that it exists

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

so do we approach it differntly?

OpenStudy (amistre64):

we can utilize the definition of a derivative: \[\lim_{x->c}\frac{g(c)-g(x)}{c-x}\]

OpenStudy (amistre64):

or \[\lim_{h->0}\frac{g(x+h)-g(x)}{h}\] \[\lim_{h->0}\frac{(x+h)^3 - (x+h)^2 + (x+h)-(x^3 - x^2 + x)}{h}\] \[\lim_{h->0}\frac{(x+h)^3 - (x^2+2xh+h^2) + x+h-x^3 + x^2 - x}{h}\] \[\lim_{h->0}\frac{(x+h)^3 - x^2-2xh-h^2 +h-x^3 + x^2 }{h}\] \[\lim_{h->0}\frac{(x+h)^3 -2xh-h^2 +h-x^3 }{h}\] \[\lim_{h->0}\frac{(x^3+3x^2h+3xh^2+h^3) -2xh-h^2 +h-x^3 }{h}\] \[\lim_{h->0}\frac{3x^2h+3xh^2+h^3 -2xh-h^2 +h }{h}\] \[\lim_{h->0}\frac{h(3x^2+3xh+h^2 -2x-h +1) }{h}\] \[\lim_{h->0}\ 3x^2+3xh+h^2 -2x-h +1\] and when h=0 we get \[3x^2 -2x +1\] but this might be a different route

OpenStudy (amistre64):

and if none of this looks familiar; then it prolly is another route that they are asking for

OpenStudy (anonymous):

how does the 2nd way taht u showed prove taht g(c) = 10

OpenStudy (anonymous):

this looks familiar..iots differnce qoutient

OpenStudy (anonymous):

but im not sure i understand how it proves taht a c exists at g(c) =10

OpenStudy (amistre64):

me either :) but then i havent got a clue as to what you are allowed to use in your proving

OpenStudy (amistre64):

all polynomials are continuous, that is a rule somewhere

OpenStudy (amistre64):

but are we trying to prove it by limits and deltas and epsilons or by some other means

OpenStudy (anonymous):

oh either way it doesnt matter...ive seen both derivitives, dffnce quotient and delta n epsilons b4

OpenStudy (amistre64):

this is "proof" but by what standards do we go by?

OpenStudy (anonymous):

we are trying to prove it by limits moslt likely

OpenStudy (anonymous):

since taht is what we have been doin in class

OpenStudy (amistre64):

the epsilon-delta version starts out like this i think g(c)-10 = e 0< |x-c| < d

OpenStudy (amistre64):

<e not =e

OpenStudy (anonymous):

how do we prove it by limits..if it is possible?

OpenStudy (amistre64):

the limit of a sum is the sum of the limits

OpenStudy (amistre64):

that might help to break it apart

OpenStudy (amistre64):

lim(x^3-x^2+x) = lim(x^3) - lim(x^2)+lim(x)

OpenStudy (amistre64):

but i got no idea how to prove it beyond htat

OpenStudy (anonymous):

oh ok but do u know how 2 do it w/ epsilon n delta since taht is how u were doing it at first

OpenStudy (amistre64):

its been to long to recall how to do it correctly; once you learn it, it tends to go by the wayside. id have to read up on it

OpenStudy (anonymous):

oh ok thanks for ur help=)

OpenStudy (amistre64):

yep, and good luck with it :)

OpenStudy (anonymous):

thanks but can u help me w/ a dffncce quotient. it confuses me b/c terhe are negative exponets

OpenStudy (amistre64):

post it up on the left and see who bites, ill get to it if i can

OpenStudy (anonymous):

ok